how might one go about factoring a larger number, such as 19,800?
Answer by mananth(12270)

19,800
...
there are many ways.
example
split the numbers with any convenient divisor.
keep factoring the split factors
19800= 198 *100
198*100=(18*11)*(10*10)
=(2*3*3)*(11)*(2*5)*(2*5)
rearrange the factors
...
m.ananth@hotmail.ca